# Tweets to Citations: Unveiling the Impact of Social Media Influencers on AI Research Visibility

### v1

@@ Abstract
As the number of accepted papers at AI and ML conferences reaches into the thousands, it has become unclear how researchers access and read research publications. In this paper, we investigate the role of social media influencers in enhancing the visibility of machine learning research, particularly the citation counts of papers they share. We have compiled a comprehensive dataset of over 8,000 papers, spanning tweets from December 2018 to October 2023, alongside 1:1 matched controls based on publication year, venue, and abstract topics. Our analysis reveals a significant increase in citations for papers endorsed by these influencers, with median citation counts 2-3 times higher than those of the control group. Additionally, the study delves into the geographic, gender, and institutional diversity of highlighted authors. These findings highlight the expanding influence of social media in scholarly communication and underscore the importance of an evolving ecosystem in today’s digital academic landscape.
@@ 1 Introduction
Acknowledging the vital function of these influencers as curators, we examine the impact of their endorsements on the citation counts of shared papers. However, we also underscore the importance of maintaining a balanced research ecosystem. An over-reliance on a select group of curators may inadvertently skew the research landscape, emphasizing certain topics or perspectives over others. Therefore, we advocate for a responsible approach to curation, encouraging influencers to maintain a journalistic integrity that includes showcasing diverse research topics, authors, and institutions.
Our study addresses the challenge of comprehending paper consumption in today’s fast-paced academic environment and seeks to unravel the factors influencing the recognition and impact of academic works amid an abundance of research and evolving norms in paper selection and review. We delve into the dual role of AK and Komatsuzaki in academic communication, observing their activity on platforms like and Hugging Face. Our primary aim is to determine whether papers endorsed by these influencers receive statistically higher citations compared to non-endorsed works, thereby shedding light on their impact on paper visibility and potential biases in the field.
@@
Comprehensive Dataset with Control Samples. We collect a dataset of shared papers from AK and Komatsuzaki containing key paper information. We select control samples through precise matching, using publication details and text embeddings as markers of paper quality. We verify that our methodology successfully controls for paper quality, evidenced by similarly distributed conference review scores. From this, we can rule out the common assumption that influencers share "higher-quality" papers, which naturally would generate more citations.
Thorough Analysis of Citations and Demographics. Our comprehensive analysis aims to determine if papers shared by these influencers receive higher citation counts compared to non-shared papers, offering insights into the impact of social media on academic research dissemination. We recognize the beneficial role of influencers in curating research amidst a deluge of publications, yet we highlight the need for balanced and diverse curation practices.
@@ 2 Related Work
Thus far, the literature has focused on social media as a forum for multi-user online discussion. In our work, we examine top-down dissemination from singular influencers, with many more followers than previously studied. Additionally, we focus on the AI/ML space, while many previous works examine fields with a much lower volume of repository publications (biology, medicine, etc.). We also contribute a geographic and gender analysis of influencer-sharing patterns and provide recommendations to conferences, influencers, and the wider community on managing the evolving field of Artificial Intelligence research.
@@ 3 Data Collection
Unique Papers | 9,171 | 1,273 |
Papers with all attributes | 8,267 | 1,194 |
Matched Papers | 7,462 | 1,152 |
@@
We model our analysis on retrospective cohort studies, in which a treatment and control group with identical underlying covariates are compared to determine the average treatment effect. In our case, we assume that a paper’s citation count is most strongly influenced by elapsed time, quality, and topic. While elapsed time is simple to measure, paper quality and topic are difficult to quantify. We use the publication venue and year as a proxy for quality and use a text embedding of the paper’s title and abstract to approximate the topic.
@@
Target Set. The Target Set is collected by searching for document identifiers in the feeds of both influencers (identifiers are found in links to arXiv.org or huggingface.co/papers). We use the Semantic Scholar API to query every document’s title, abstract, year of publication, venue of publication, and citation count. We also collect the authors of each paper but do not use this information in our matching. Notably, some papers are either not available on Semantic Scholar or are not freely accessible. Hence, we remove any paper that lacks a required attribute. The number of unique papers and papers with all attributes for each influencer is shown in Table 1.
@@
Control Set. To build our Control Set, we first collect a large dataset of papers presented at the same venues and in the same years as those in the Target Set. Specifically, for every instance of a paper published in year at venue , we query the Semantic Scholar API for all papers published in year at venue . We sample as many papers as the Semantic Scholar Database has available, yielding 247,993 unique papers, and 124,940 papers with all required attributes. This dataset forms the corpus from which we match papers to our target set.
@@
The choice of matching algorithm has been the subject of fierce debate; we follow the recommendations in King and Nielsen (2019) by performing exact matching on our categorical variables (publication venue and topic) and Euclidean distance matching on our continuous variables (topic embedding). Note that Euclidean distance is equivalent to cosine similarity for normalized vectors. Optimal matching can be reduced to the linear sum assignment problem, for which many efficient algorithms exist Crouse (2016). We use the implementation available in SciPy Virtanen et al. (2020). Alternate matching algorithms did not significantly change our results.
@@
Using this methodology, we match each paper in the Target Set to a paper in the Control Set with respect to the gathered covariates. We exclude any paper for which we cannot find an exact match on publication venue or year. For matching on text embeddings, we set a cutoff of 0.6 for cosine similarity, which ensures close topic similarity in matched pairs while retaining 91% of AK’s tweeted papers and 96% of Komatsuzaki’s tweeted papers.
Qualitatively, the matched pairs are very similar in topic, almost always covering the same subfield of research (for example, diffusion models applied to image generation), approaching the same problems, and using similar or the same methods. We note that the modal cosine similarity is 0.7, where performance is much better than at the cutoff value of 0.6, but we show matched pairs across a range of cosine similarities to demonstrate that our method is successful at all levels.
As before, we express caution about dropping papers from the Target Set, since this may introduce bias. However, we find similar results when we do not drop any papers from the Target Set. For more details on dropping papers, alternate matching algorithms, and example matched pairs, see Appendix A.
@@
Review Scores. To ensure that we have successfully controlled for quality through our matching, we will look at the review scores of experimental and control pairs from some selected conferences. We extract the review data from ICLR, AAAI, ICML, NeurIPS, EMNLP, and KDD using OpenReview (2023). Across both paired datasets, we found 991 out of 8525 unique pairs with available scores.
@@
We plot the mean review score of each treatment paper against the mean score of its paired control (Figure 2). For conferences that do not use numbered review scores, we assign numbers based on those of other conferences (7: Accept, 5: Borderline Accept, 3: Borderline Reject, 2: Reject). Using the same three significance tests from Table 3, we do not find sufficient evidence to reject the null hypothesis that the control and experimental scores are from the same distribution (-value ). Assuming that mean review scores are an accurate measure of paper quality, we can conclude that we have effectively controlled for paper quality in our datasets.
@@ 4 Analysis
The Q-Q results are further strengthened by the Cliff’s Delta values for each pairwise sample (Table 3). Using the effect sizes proposed by Hess and Kromrey, the Cliff’s Delta shows a large and medium effect for Komatsuzaki’s and AK’s sharing, respectively (2004). This suggests that the impact of influencer sharing is practically significant in altering the outcome variable, which is the citation count of the papers.
Finally, we establish statistical significance with three tests comparing the distributions of the experimental data with that of the control sets, Epps-Singleton (ES), Kolmogorov-Smirnov (KS), and Mann-Whitney U (MWU), none of which assume normal distribution, which is essential for our data (especially controls). Table 3 shows the results, all with -values well below even a stringent . From this, we can strongly reject the null hypothesis that the citation distributions for the influencer-shared papers and the control papers are the same.
AK | Komatsuzaki | |
@@
Mann-Whitney U | ||
Cliff’s Delta | 0.227 | 0.333 |
@@
To test this, we first collect the geographic data of the shared papers. First, we use Semantic Scholar and dblp (2023) to collect the affiliation data of all listed authors from each test set. We then utilize the Nominatim geocoding API to find the approximate latitude and longitude of each affiliation. However, many of the reported affiliations are not suitable for the API and produce incorrect geocodes. To combat this, we manually adjust visibly inaccurate coordinates using publically available addresses online. From this information, we perform a reverse geocode with Nominatim to find the country of each affiliation, then use majority voting to assign each publication a country. At this point, we can see that both influencers share papers from around the world in Figure 5. Finally, we aggregate these countries into the same geographic areas used in the HAI Report and plot using a similar format (Figure 7).
To account for the discrepancy between date ranges in the HAI Report and our influencers’ activity, we will limit our analysis to the overlap between them. Additionally, we will focus on the range from 2018 to 2021, because of the low sample size of shared papers pre-2018–only 1 for AK and 4 for Komatsuzaki.
During these years, Figure 6 indicates a slight decline in the United States’ share of AI repository publications following its peak. This decrease may suggest a maturation of the AI field, where research is becoming more globally dispersed. Concurrently, the European Union and United Kingdom demonstrate a modest uptick in publications after a consistent decline from 2010-2017, while China’s share continues to rise.
In contrast, the sharing patterns of the influencers from 2018 to 2021, shown in Figure 7, demonstrate a notable deviation from these global trends. Specifically, AK’s shared publications exhibit a sharp decline in the initially overwhelming "Unknown" category, with a concurrent and dramatic rise in the United States’ share. This seems to indicate better affiliation reporting rather than a change in AK’s sharing practices, as the shares from other areas stay relatively constant. Komatsuzaki’s data exhibits a consistent focus on U.S.-affiliated papers, with the notable absence of some geographic regions until later years.
While the global landscape of AI publications suggests an increase in diversity and a more even distribution of research output, our data presents a skewed alignment, favoring the United States. We must note, though, that using solely self-reported affiliations can have an inherent bias toward the United States. For example, many researchers affiliated with multinational organizations are assigned to the United States (where headquarters are located) despite working with a branch in another area. Additionally, we must note the prominence of the "Unknown" category in the data of both influencers, where affiliations were not found.
@@ 5 Discussion
Influencers and the ML Community. Influencers serve as pivotal curators in the ML landscape. With the explosive growth of machine learning research, the community increasingly relies on social media to keep up-to-date with new developments. This reliance has naturally led to the rise of centralized influencers. Their role in streamlining the dissemination process is akin to that of journalists in news media, making novel ideas and breakthroughs more accessible. Just as journalists are indispensable to civic society, influencers enable the ML community to stay informed on the bleeding edge.
However, we urge caution against excessive dependence on a limited set of information sources. Research inherently involves bottom-up exploration of a wide array of topics to uncover groundbreaking ideas. Focusing on a handful of individuals’ highlighted papers necessarily offers a narrow view of the research landscape. We encourage the community to keep the online academic space competitive, allowing for a diversity of highly visible ideas. This goal can be achieved through active participation in an open, community-driven curation process, enhancing the variety of prominent ideas.
Additionally, like how we expect journalists to maintain integrity–through highlighting both prominent and underreported stories–influencers have a responsibility to share a diverse array of ideas. This includes showcasing various techniques and subtopics within their areas of focus, thereby exposing community members to new approaches and concepts.
@@
While we specifically examine geography and gender, these principles apply broadly: proactive efforts in the online domain can significantly enhance equity. We encourage influencers to actively promote visibility for diverse groups globally.
The geographical and social background of an influencer could inadvertently influence the diversity of shared papers. We emphasize that this does not necessarily indicate bias among current influencers but serves as a call for vigilance against potential biases as the AI/ML domain expands. This connects back to our recommendation for fostering a diverse online space where everyone can contribute to broadening the spectrum of shared ideas.
@@ 6 Conclusion
In this study, we have explored the evolving landscape of academic discourse in AI/ML, focusing on the role of social media influencers in the dissemination and recognition of scholarly works. Our investigation into the activities of AK and Aran Komatsuzaki on and Hugging Face has provided crucial insights into how influencers are shaping the visibility and impact of research papers in this rapidly advancing field.
Influencers as Catalysts for Visibility: Our comprehensive analysis reveals that papers shared by AK and Komatsuzaki receive statistically higher citation counts compared to non-endorsed works, confirming the significant role these influencers play in amplifying the reach of specific research. This influence is not merely a reflection of sharing ’higher-quality’ papers, as controlled by our dataset, but a testament to their ability to highlight and contextualize substantial findings for the community.
Influencers as Curators: Influencers serve a key role in the AI/ML space by making novel ideas and breakthroughs more accessible. However, it’s crucial to balance their influence with a commitment to diversity. The reliance on a select few for information could result in a limited perspective on the vast research landscape. Therefore, we advocate for a diverse, competitive online academic space where a wide array of ideas and research can gain visibility.
Implications for Academic Practices: Our findings have significant implications for how the AI/ML academic community approaches the dissemination and reception of research. The rising influence of social media figures in academic communication necessitates a reevaluation of traditional modes of paper selection and review. We propose that conference organizers and academic institutions engage in a dialogue to evolve the conference system and peer-review processes, ensuring they adapt to the changing norms and continue to recognize and disseminate quality research effectively.
Future Directions: This study opens several avenues for future research. One potential area is to investigate the extent to which the trends observed in AI/ML apply to other scientific domains. Additionally, further exploration into the mechanisms behind the influence of social media on academic recognition, such as the role of algorithms and network effects, could provide deeper insights into managing the balance between traditional and modern dissemination channels.
In conclusion, our research highlights the growing interplay between social media and academic research dissemination in the AI/ML community. The influencer phenomenon, while beneficial in curating and highlighting significant research, brings with it challenges that necessitate careful consideration and action from the academic community. By fostering responsible curation practices and evolving traditional academic systems, we can ensure a balanced and inclusive research ecosystem that recognizes and promotes a wide array of scholarly contributions.

### v2

@@ Abstract
As the number of accepted papers at AI and ML conferences reaches into the thousands, it has become unclear how researchers access and read research publications. In this paper, we investigate the role of social media influencers in enhancing the visibility of machine learning research, particularly the citation counts of papers they share. We have compiled a comprehensive dataset of over 8,000 papers, spanning tweets from December 2018 to October 2023, alongside controls precisely matched by 9 key covariates. Our statistical and causal inference analysis reveals a significant increase in citations for papers endorsed by these influencers, with median citation counts 2-3 times higher than those of the control group. Additionally, the study delves into the geographic, gender, and institutional diversity of highlighted authors. Given these findings, we advocate for a responsible approach to curation, encouraging influencers to uphold the journalistic standard that includes showcasing diverse research topics, authors, and institutions.
@@ 1 Introduction
Acknowledging the vital function of these influencers as curators, we examine the impact of their endorsements on the citation counts of shared papers. However, we also underscore the importance of maintaining a balanced research ecosystem. An over-reliance on a select group of curators may inadvertently skew the research landscape, emphasizing certain topics or perspectives over others. Therefore, we advocate for a responsible approach to curation, encouraging influencers to maintain the journalistic standard that includes showcasing diverse research topics, authors, and institutions.
@@
Comprehensive Dataset with Control Samples. We collect a dataset of shared papers from AK and Komatsuzaki containing key paper information. We select control samples through precise matching, using several paper and author covariates.
Thorough Analysis of Citations and Demographics. Our comprehensive analysis aims to determine if papers shared by these influencers receive higher citation counts compared to non-shared papers. We conduct statistical and causal inference analysis and find significant evidence for a direct effect. We then explore the geographic and gender distributions of influencer-shared papers and compare them to reference points from the community at large.
@@ 2 Related Work
We note that the AI/ML community is much more active on platforms like . While many previous works examine fields with a much lower volume of repository publications, the short research cycle and the need for quick dissemination of results, coupled with the rise of arXiv, motivates us to study the problem from the perspectives of modern AI/ML research lens. Additionally, the literature thus far has focused on social media as a forum for multi-user online discussion. In our work, we examine top-down dissemination from singular influencers, with many more followers than previously studied. We also contribute a geographic and gender analysis of influencer-sharing patterns and provide recommendations to conferences, influencers, and the wider community on managing the evolving field of Artificial Intelligence research.
@@ 3 Data Collection
# Unique Papers | 9,171 | 1,273 |
# with all attributes | 8,259 (90%) | 1,191 (94%) |
# Matched Papers | 6,890 (75%) | 955 (78%) |
@@
We model our analysis on retrospective cohort studies, in which a treatment and control group with identical underlying covariates are compared to determine the average treatment effect. In our case, we assume that a paper’s citation count is most strongly influenced by elapsed time, quality, topic, and author prominence. While elapsed time is simple to measure, paper quality, topic are difficult to quantify. We use the publication venue and year as a proxy for quality and use a text embedding of the paper’s title and abstract to approximate the topic. We use first author total citation counts and h-index, and the maximum of all authors’ citation counts and h-indexes to adjust for their prominence.
@@
Target Set. The Target Set is collected by searching for document identifiers in the feeds of both influencers (identifiers are found in links to arXiv.org or huggingface.co/papers). We use the Semantic Scholar (S2) API to query every document for desired attributes (Table 4). Notably, some papers are either not available on S2 or are not freely accessible. Hence, we remove any paper that lacks any required attribute. The number of unique papers and papers with all attributes for each influencer is shown in Table 2.
@@
title, abstract, year, venue, |
citationCount, |
embedding.specter_v2, |
authors.{name, affiliations, |
hIndex, citationCount} |
year, venue, |
open_access, n_authors, |
first(h-index), max(h-index), |
first(citations), max(citations), |
topic_embedding |
@@
Control Set. To build our Control Set, we first collect a large dataset of papers presented at the same venues and in the same years as those in the Target Set. Specifically, for every instance of a paper published in year at venue , we query S2 for all papers published in year at venue .
@@
We follow the recommendations in King and Nielsen (2019) by performing exact matching on our categorical (year, venue, open_access111Due to the nature of influencer sharing, all sampled papers are open access.) and binned variables (n_authors, h-index, author cites), and Euclidean distance222Note that Euclidean distance is equivalent to cosine similarity for normalized vectors. matching on our continuous variables (topic_embedding). Optimal matching can be reduced to the linear sum assignment problem, for which many efficient algorithms exist Crouse (2016). We use the implementation available in SciPy Virtanen et al. (2020).
@@
Using this methodology, we match each paper in the Target Set to a paper in the Control Set with respect to the gathered covariates (Table 4). We exclude any paper for which we cannot find an exact match on categorical variables. For matching on topic embeddings, we use SPECTER2 Singh et al. (2022). Matches between papers were strong, and the distribution of cosine similarities as well as example matched pairs are detailed in Appendix A.
@@
Review Scores. To determine if we have successfully controlled for quality through our matching, we will look at the review scores of experimental and control pairs from some selected conferences. We extract the review data using OpenReview (2023). Across both paired datasets, we found 939 out of 7222 unique pairs with available scores.
@@
We plot the mean review score of each treatment paper against the mean score of its paired control (Figure 2). For conferences that do not use numbered review scores, we assign numbers based on those of other conferences (7: Accept, 5: Borderline Accept, 3: Borderline Reject, 2: Reject).
Using the same three significance tests from Table 5, we do not find sufficient evidence to reject the null hypothesis that the control and experimental scores are from the same distribution (-value ). Assuming that mean review scores are an accurate measure of paper quality, we conclude that we have effectively controlled for paper quality in our matching.
@@ 4 Analysis
Test | AK | Komatsuzaki |
@@
Mann-Whitney U |
Finally, we establish statistical significance with three tests comparing the distributions of the experimental data with that of the control sets, Epps-Singleton (ES), Kolmogorov-Smirnov (KS), and Mann-Whitney U (MWU), none of which assume normal distribution, which is essential for our data. Table 5 shows the results, all with -values well below even a stringent . From this, we can strongly reject the null hypothesis that the citation distributions for the influencer-shared papers and the control papers are the same.
@@
Causal Inference: Setup. Although we have shown a significant correlation between influencer sharing and citations, we have not investigated a causal link. For this, we turn to the model, techniques, and assumptions presented in Elazar et al. (2023). For additional background and details, see Appendix C
We aim to estimate the causal effect of the treatment , indicating that the paper was shared by influencers333To increase sample size and simplify our analysis, we combine our test sets for this portion of the analysis., on the outcome , if the paper is "highly-cited" or "less-cited." Besides , there are a number of other factors, known as confounders, that can affect . We divide these confounders into two sets: observed –which are the same used in our matching (Section 3)–and unobserved –which are difficult to quantify or measure (e.g. novelty, contribution, hype).
To debias the effects of unobserved confounders , we employ the negative outcome control (NOC) framework Card and Krueger (1993); Lipsitch et al. (2010). By finding a negative control outcome that shares the same confounders as , but is not causally affected by , we can attempt to correct for the bias introduced by . To this end, we select a paper’s average review score for , which we have already shown is not significantly correlated to (Figure 2) and which we believe share many of the same confounders (e.g., quality, contribution, author experience). However, another NCO can be substituted if deemed more suitable.
To account for fine-grained citation counts and review scores in comparison to the binary nature of and , we define and . For both, their value is 1 if the corresponding real-valued quality (citation count or average review score) is above the (or ) quantile of the relevant sample distribution. We will use values of for both and in our analysis.
Finally, we estimate the causal effect of influencer sharing on paper citations with the average treatment effect of the treated. Using the difference-in-difference assumption with NOC, we can write our estimate Elazar et al. (2023):
where ATET values closer to indicate weak to no effect of the treatment, and large absolute values indicate a stronger effect.
Causal Inference: Results. First, we estimate the effect of on without accounting for the effects of . We estimate ATET (95%-bootstrapping CI) to be 19% (15-24), 16% (13-20), 9% (6-12) for , respectively. This confirms the significant association between influencer sharing and citations we investigated during the contrasting analysis. However, this does not preclude the existence of strong confounders.
For that, we will use NOC to attempt to debias the effects of unobserved confounders, including quality. We record our results in Figure 5 and observe the following: (1) debiased effects are smaller than the unadjusted estimates, indicating our NCO has succeeded in accounting for some unobserved confounders; (2) all of our estimated effects are significant to the 95% level, because none contain 0 in their confidence interval; (3) influencer sharing is most successful at increasing the citations in lower percentiles–papers with very high citation counts experience a smaller (yet still significant) effect compared to those with fewer citations.
@@
To test this, we first collect the geographic data of the shared papers. First, we use S2 and dblp (2023) to collect the affiliation data of all listed authors from each test set. We then utilize the Nominatim geocoding API to find the approximate latitude and longitude of each affiliation, manually adjusting visibly inaccurate coordinates. From this information, we find the country of each affiliation and then use majority voting to assign each publication a geographic area. At this point, we can see that both influencers share papers from around the world in Figure 6. Finally, we aggregate these countries into the same geographic areas used in the HAI Report and plot using a similar format (Figure 8).
To account for the discrepancy between date ranges in the HAI Report and our influencers’ activity, we will limit our analysis to the overlap between them. Additionally, we will focus on the range from 2018 to 2021, because of the low sample size of shared papers pre-2018–only 5 in total.
During these years, Figure 7 indicates a slight decline in the United States’ share of AI repository publications following its peak. Concurrently, the European Union and United Kingdom demonstrate a modest uptick in publications after a consistent decline from 2010-2017, while China’s share continues to rise.
In contrast, the sharing patterns of the influencers from 2018 to 2021, shown in Figure 8, demonstrate a notable deviation from these global trends. Specifically, both influencers’ data exhibits a consistent focus on U.S.-affiliated papers, with a much smaller, relatively constant portion from other areas. The increase in U.S.-affiliated papers seems to be a change in reporting rather than a change in sharing practices.
We must note, though, that using solely self-reported affiliations can have an inherent bias toward the United States. For example, many researchers affiliated with U.S.-based organizations are assigned to the United States despite working out of another area. Additionally, we must note the prominence of the "Unknown" category in the data of both influencers, where affiliations were not found.
@@ 5 Discussion
Influencers and the ML Community. Influencers serve as pivotal curators in the ML landscape. With the explosive growth of machine learning research, the community increasingly relies on social media to keep up-to-date with new developments. Their role in streamlining the dissemination process is akin to that of journalists in news media, making novel ideas and breakthroughs more accessible.
However, we advise caution against excessive dependence on a limited set of information sources. Research inherently involves bottom-up exploration of a wide array of topics. Focusing on a handful of individuals’ highlighted papers necessarily offers a narrow view of the research landscape. We encourage the community to keep the online academic space competitive, allowing for a diversity of highly visible ideas. This goal can be achieved through active participation in an open, community-driven curation process, enhancing the variety of prominent ideas. Additionally, we urge influencers to share a diverse array of ideas. This includes showcasing various techniques and subtopics within their areas of focus, thereby exposing community members to new approaches and concepts.
@@
While we specifically examine geography and gender, these principles apply broadly: proactive efforts in the online domain can significantly enhance equity. We encourage influencers to actively promote visibility for diverse groups globally. We emphasize that this does not necessarily indicate bias among current influencers but serves as a call for vigilance against potential biases as the AI/ML domain expands. This connects back to our recommendation for fostering a diverse online space where everyone can contribute to broadening the spectrum of shared ideas.
@@ 6 Conclusion
This study delves into the influence of social media influencers on the dissemination and recognition of academic work in AI/ML, specifically examining the impact of AK and Aran Komatsuzaki on platforms like and Hugging Face. Our analysis demonstrates that papers endorsed by these influencers receive significantly more citations than those that are not, underscoring the role of influencers in not only amplifying the reach of specific research but also in shaping its visibility in the field. This influence extends beyond merely sharing higher-quality papers, highlighting their ability to effectively contextualize and promote substantial findings within the community.
The study further discusses the dual role of influencers as both catalysts for visibility and curators of content, emphasizing the need for a balance in their influence to ensure a diversity of perspectives in the AI/ML research landscape. The implications of this influence are profound, suggesting a need for the academic community to reassess traditional methods of paper selection and review. Our findings encourage a dialogue among conference organizers and academic institutions to evolve the peer-review process and adapt to these changing norms. Additionally, the study opens avenues for future research into the generalizability of these trends in other scientific fields and the mechanisms behind the influence of social media on academic recognition.
Impact Statement
Recognizing the influential role of social media influencers in the academic community, particularly in the fields of AI/ML, it’s important to gently encourage them to be aware of their impact. Influencers play a key role in shaping discussions and trends, and there is a growing need for them to consider the diversity and inclusivity of the research they share. This is not to undermine their contributions but to enhance the richness of academic discourse. Similarly, the academic community might benefit from reevaluating traditional metrics of research impact, embracing both peer-reviewed channels and digital platforms for a more holistic approach to recognizing scholarly work.
In doing so, we hope to spark future research into understanding any unintentional biases influencers might have, such as favoring research from prestigious institutions or specific regions. This is a subtle yet crucial aspect of ensuring a balanced representation in the academic landscape, preventing the overshadowing of significant but less publicized work. While influencers may have personal or commercial interests, we trust in their capacity to acknowledge and manage these to avoid conflicts of interest. This approach would help in preventing echo chambers and promoting a diverse range of perspectives in AI/ML research. Overall, we advocate for a collaborative and mindful effort among influencers and the academic community to foster an equitable and inclusive environment for scholarly discussions.